# Compositional Design of Safety Controllers for Large-scale Stochastic Hybrid Systems

### v1

@@ 4. Compositional Construction of A-CBC
    Flow case: $10\leq\vartheta_{i}\leq\varepsilon_{2_{i}}-$ 1, $x_{i}^{\prime}=x_{x_{i}\nu_{i}w_{i}}(\tau^{-})$, and $\vartheta_{i}^{\prime}=\vartheta_{i}+1$;
@@
Consider an interconnected A-SHS $\mathcal{I}(\mathcal{A}_{1}(\Phi_{1}),\ldots,\mathcal{A}_{N}(\Phi_{N}))$ induced by $N\in\mathbb{N}^{+}$ A-SHSs $\mathcal{A}_{i}(\Phi_{i})$. Suppose that each $\mathcal{A}_{i}(\Phi_{i})$ admits an A-CSBC $\mathcal{B}_{i}$, as defined in Definition 3.1. For a vector $\xi$ with $\xi_{i}>0,\forall i\in\{1,\dots,N\}$, if
@@
$\beta:=\sum_{i=1}^{N}\xi_{i}\beta_{i},\quad\mu:=\sum_{i=1}^{N}\xi_{i}\mu_{i},\quad\eta:=\sum_{i=1}^{N}\xi_{i}\eta_{i},\quad\gamma:=1+\Gamma,~{}\text{with}~{}\max_{1\leq i\leq N}\Gamma_{i}<\Gamma<0,~{}\text{and}~{}\Gamma\in(-1,0).$
@@
$\gamma s$ $:=\max\Big{\{}s+\xi^{\top}(-\Theta+\Psi)\tilde{\mathcal{B}}(x,\vartheta)\,\big{|}\,\xi^{\top}\tilde{\mathcal{B}}(x,\vartheta)=s\Big{\}},\quad\eta:=\sum_{i=1}^{N}\xi_{i}\eta_{i},$
with ~$\tilde{\mathcal{B}}(x,\vartheta)=\big{[}\mathcal{B}_{1}(x_{1},\vartheta_{1});\ldots;\mathcal{B}_{N}(x_{N},\vartheta_{N})\big{]}$, condition (3.2c) is also met. The only remaining part is to show that $\gamma=1+\Gamma$ and $10<\gamma<$ 1. Given that $\xi^{\top}(-\Theta+\Psi):=\big{[}\xi_{1}\Gamma_{1};\dots;\xi_{N}\Gamma_{N}\big{]}^{\top}<0$, and $\max_{1\leq i\leq N}\Gamma_{i}<\Gamma<0$ with $\Gamma\in(-1,0)$, we have
@@
$=s+\big{[}\xi_{1}\Gamma_{1};\dots;\xi_{N}\Gamma_{N}\big{]}^{\top}\big{[}\mathcal{B}_{1}(x_{1},\vartheta_{1});\ldots;\mathcal{B}_{N}(x_{N},\vartheta_{N})\big{]}$
@@
As demonstrated in Theorem 4.2, under the compositionality conditions of (4.1a)-(4.1b), an A-CBC for the interconnected A-SHS can be compositionally constructed based on A-CSBCs of individual subsystems. In the next section, we provide the necessary conditions for constructing the A-CSBC for the augmented SHS.
5. Construction of A-CSBC
In this section, we establish conditions on the SHS $\Phi_{i}$ that allow for the construction of an A-CSBC for $\mathcal{A}_{i}(\Phi_{i})$. To do so, we first define a notion of control barrier certificates (CBC) over original $\Phi_{i}$.
@@ 5. Construction of A-CSBC
Given an SHS $\Phi_{i}$ with initial and unsafe sets $X_{0_{i}},X_{u_{i}}\subseteq X_{i}$, a function ¯$\bar{\mathcal{B}}_{i}:X_{i}\rightarrow\mathbb{R}_{0}^{+}$ is called a control barrier certificate (CBC) for $\Phi_{i}$ if there exist $\gamma_{1_{i}}\in{\mathbb{R}},\bar{\kappa}_{i},\gamma_{2_{i}}\in{\mathbb{R}}^{+}$, $\bar{\mu}_{i},\bar{\beta}_{i},\bar{\varphi}_{i_{1}},\bar{\varphi}_{i_{2}},\bar{\eta}_{1_{i}},\bar{\eta}_{2_{i}}\in{\mathbb{R}}_{0}^{+}$, such that
¯$\bar{\mathcal{B}}_{i}(x_{i})\geq\bar{\kappa}_{i}\|x_{i}\|^{2},\quad\quad\quad\quad~{}~{}~{}~{}~{}\forall x_{i}\in X_{i},$ (5.1a)
¯$\bar{\mathcal{B}}_{i}(x_{i})\leq\bar{\mu}_{i},\quad\quad\quad\quad\quad\quad\!\!\forall x_{i}\in X_{0_{i}},$ (5.1b)
¯$\bar{\mathcal{B}}_{i}(x_{i})\geq\bar{\beta}_{i},\quad\quad\quad\quad\quad\quad\!\forall x_{i}\in X_{u_{i}},$ (5.1c)
@@
¯$\mathcal{L}\bar{\mathcal{B}}_{i}(x_{i})\leq-\gamma_{1_{i}}\bar{\mathcal{B}}_{i}(x_{i})+\bar{\varphi}_{i_{1}}\|w_{i}\|^{2}+\bar{\eta}_{1_{i}},$ (5.1d)
where ¯$\mathcal{L}\bar{\mathcal{B}}_{i}(x_{i})$ is the infinitesimal generator acting on ¯$\bar{\mathcal{B}}_{i}(x_{i})$, defined as [Oks13]
¯$\mathcal{L}\bar{\mathcal{B}}_{i}(x_{i})$ =¯$=\frac{\partial\bar{\mathcal{B}}_{i}(x_{i})}{\partial x_{i}}\!f_{1_{i}}\!(x_{i},\!\nu_{i},\!w_{i})+\frac{1}{2}\!\textsf{Tr}~{}\!(\sigma_{i}(x_{i})\sigma_{i}(x_{i})\!^{\top}\frac{\partial^{2}{\bar{\mathcal{B}}_{i}(x_{i})}}{\partial x_{i}\partial x_{j}})+\sum_{j=1}^{\textsf{r}_{i}}\lambda_{j_{i}}(\,\bar{\mathcal{B}}_{i}(x_{i}+\rho_{i}(x_{i})\textsf{e}_{j_{i}}^{\textsf{r}_{i}})-\bar{\mathcal{B}}_{i}(x_{i})),$
where $\lambda_{j}$ is the rate of Poisson processes and $\textsf{e}_{j}^{\textsf{r}}$ is an r-dimensional vector with $1$ on the $j$-th entry and 00 elsewhere;
@@
Before offering the main result of this section, we also raise the following lemma to quantify an upper bound on the evolution of CBC ¯$\bar{\mathcal{B}}_{i}$.
@@
$\mathds{E}$ [¯$\Big{[}\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t_{k+1}^{-}))\,\big{|}\,x_{i},\nu_{i},w_{i}\Big{]}\leq e^{-\gamma_{1_{i}}(t_{k+1}-t_{k})}\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t_{k})\!)+e^{-\gamma_{1_{i}}(t_{k+1}-t_{k})}(t_{k+1}\!-\!t_{k})(\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+\bar{\eta}_{1_{i}}).$ (5.2)
Proof. From Dynkin’s formula [Dyn65], we have:
$\mathds{E}$ [¯$\Big{[}\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t_{k+1}^{-}))\,\big{|}\,x_{i},\nu_{i},w_{i}\Big{]}=\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t_{k}))\!+\!\mathds{E}\Big{[}\int_{t_{k}}^{t_{k+1}}\!\!\!\!\!\!\!\mathcal{L}\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t))\mathsf{d}t\Big{]}\!.$
@@
$\mathds{E}\Big{[}\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t_{k+1}^{-}))\,\big{|}\,x_{i},\nu_{i},w_{i}\Big{]}$$\leq\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t_{k}))+\mathds{E}\Big{[}\int_{t_{k}}^{t_{k+1}}\big{(}-\gamma_{1_{i}}\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t))+\bar{\varphi}_{i_{1}}\|w_{i}\|^{2}+\bar{\eta}_{1_{i}}\big{)}\mathsf{d}t$
=¯$=\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t_{k}))\!+\!(t_{k+1}\!-\!t_{k})(\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+\bar{\eta}_{1_{i}})-\int_{t_{k}}^{t_{k+1}}\!\!\gamma_{1_{i}}\mathds{E}\Big{[}\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t))\Big{]}\mathsf{d}t.$
By employing Gronwall inequality [Gro19], we have
$\mathds{E}$ [¯$\Big{[}\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t_{k+1}^{-}))\,\big{|}\,x_{i},\nu_{i},w_{i}\Big{]}\leq e^{-\gamma_{1_{i}}(t_{k+1}-t_{k})}\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t_{k}))+e^{-\gamma_{1_{i}}(t_{k+1}-t_{k})}(t_{k+1}\!-\!t_{k})(\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+\bar{\eta}_{1_{i}}),$
@@
Under Definition 5.1 and Lemma 5.2, we propose the next theorem to construct an A-CSBC for $\mathcal{A}_{i}(\Phi_{i})$ based on the CBC ¯$\bar{\mathcal{B}}_{i}$.
@@
Given an SHS $\Phi_{i}=(X_{i},U_{i},\mathcal{U}_{i},W_{i},\mathcal{W}_{i},$ $\sigma_{i},\rho_{i},f_{1_{i}},\varsigma_{i},f_{2_{i}})$ with its associated A-SHS $\mathcal{A}_{i}(\Phi_{i})=(\mathbb{X}_{i},\mathbb{U}_{i},\mathds{W}_{i},\sigma_{i},\rho_{i},\varsigma_{i},\mathbb{F}_{i},\mathbb{Y}_{i},\mathbb{H}_{i})$, suppose ¯$\bar{\mathcal{B}}_{i}$ is a CBC for $\Phi_{i}$. If
@@
$\varpi_{i}:=\left\{\begin{array}[]{lr}1,\quad\quad\quad\quad\,~{}~{}~{}~{}\text{if}~{}\gamma_{1_{i}}>0~{}\&~{}0<\gamma_{2_{i}}<1,\\ e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\vartheta_{i}},\quad\text{if}~{}\gamma_{1_{i}}>0~{}\&~{}\gamma_{2_{i}}\geq 1,\\ \gamma_{2_{i}}^{\frac{\vartheta_{i}}{\alpha_{2_{i}}}},\quad\quad\quad\quad~{}~{}~{}~{}\!\!\text{if}~{}\gamma_{1_{i}}\leq 0~{}\&~{}0<\gamma_{2_{i}}<1,\end{array}\right.$ (5.8)
for some $10<\alpha_{1_{i}}<$ 1 and $\alpha_{2_{i}}>\varepsilon_{2_{i}}$.
@@
$\mathds{E}\Big{[}\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t_{k+1}^{-}))\,\big{|}\,x_{i},\nu_{i},w_{i}\Big{]}$ $\leq e^{-\gamma_{1_{i}}\tau}\big{(}\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(0))\!+\!\tau(\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+\bar{\eta}_{1_{i}})\big{)}$
@@
      $=e^{-\gamma_{1_{i}}\tau}\mathcal{B}_{i}(x_{i},\!\vartheta_{i})\!+\!e^{-\gamma_{1_{i}}\!\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+e^{-\gamma_{1_{i}}\tau}\!\tau\bar{\eta}_{1_{i}}$
      $\leq e^{-\gamma_{1_{i}}\tau}\mathcal{B}_{i}(x_{i},\!\vartheta_{i})\!+\!e^{-\gamma_{1_{i}}\!\tau}\tau\bar{\varphi}_{i_{1}}(\|w_{i}\|_{\infty}^{2}\!+\!\|w_{i}\|^{2})+e^{-\gamma_{1_{i}}\tau}\!\tau\bar{\eta}_{1_{i}}$
      $=e^{-\gamma_{1_{i}}\tau}\mathcal{B}_{i}(x_{i},\!\vartheta_{i})\!+\!e^{-\gamma_{1_{i}}\!\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|^{2}+e^{-\gamma_{1_{i}}\!\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+e^{-\gamma_{1_{i}}\tau}\!\tau\bar{\eta}_{1_{i}}.$
@@
      $\leq\gamma_{2_{i}}\bar{\mathcal{B}}_{i}(x_{i})\!+\!\bar{\varphi}_{i_{1}}\|w_{i}\|^{2}\!+\!\bar{\eta}_{2_{i}}=\gamma_{2_{i}}\mathcal{B}_{i}(x_{i},\vartheta_{i})\!+\!\bar{\varphi}_{i_{2}}\|w_{i}\|^{2}+\bar{\eta}_{2_{i}}.$
      By defining $\gamma_{i}=\max\{e^{-\gamma_{1_{i}}\tau},\gamma_{2_{i}}\}$, $\varphi_{i}=\max\{e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}},\bar{\varphi}_{i_{2}}\}$ and $\eta_{i}\!=\!\max\{e^{-\gamma_{1_{i}}\!\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+e^{-\gamma_{1_{i}}\tau}\!\tau\bar{\eta}_{1_{i}},\bar{\eta}_{2_{i}}\}$, we have
      $\mathds{E}\Big{[}\mathcal{B}_{i}(x_{i}^{\prime},\!\vartheta_{i}^{\prime})\big{|}x_{i},\!\nu_{i},\!w_{i},\!\vartheta_{i}\Big{]}$ $\leq\gamma_{i}\mathcal{B}_{i}(x_{i},\!\vartheta_{i})+\varphi_{i}\|w_{i}\|^{2}\!+\!\eta_{i}.$
@@
      $\hskip 56.9055pt\mathds{E}$ $\Big{[}\mathcal{B}_{i}(x_{i}^{\prime},\vartheta_{i}^{\prime})\,\big{|}\,x_{i},\nu_{i},w_{i},\vartheta_{i}\Big{]}$

### v2

@@ 4 Compositional Construction of A-CBC
    Flow case: $0\leq\vartheta_{i}\leq\varepsilon_{2_{i}}-1$, $x_{i}^{\prime}=x_{x_{i}\nu_{i}w_{i}}(\tau^{-})$, and $\vartheta_{i}^{\prime}=\vartheta_{i}+1$;
@@
Consider an interconnected A-SHS $\mathcal{I}(\mathcal{A}_{1}(\Phi_{1}),\ldots,\mathcal{A}_{N}(\Phi_{N}))$ induced by $N\in\mathbb{N}^{+}$ A-SHSs $\mathcal{A}_{i}(\Phi_{i})$. Suppose that each $\mathcal{A}_{i}(\Phi_{i})$ admits an A-CSBC $\mathcal{B}_{i}$, as defined in Definition 3.1. For a vector $\xi$ with $\xi_{i}>0$, and a scalar value $\Gamma\in(-1,0)$, where $\max_{1\leq i\leq N}\Gamma_{i}<\Gamma<0$, $\forall i\in\{1,\dots,N\}$, if
@@
$\beta:=\sum_{i=1}^{N}\xi_{i}\beta_{i},\;\mu:=\sum_{i=1}^{N}\xi_{i}\mu_{i},\;\eta:=\sum_{i=1}^{N}\xi_{i}\eta_{i},\;\gamma:=1+\Gamma.$
@@
$\gamma s$ $:=\max\Big\{s+\xi^{\top}(-\Theta+\Psi)\tilde{\mathcal{B}}(x,\vartheta)\,\big|\,\xi^{\top}\tilde{\mathcal{B}}(x,\vartheta)=s\Big\},$
$\eta$ $:=\sum_{i=1}^{N}\xi_{i}\eta_{i},$
with $\tilde{\mathcal{B}}(x,\vartheta)=\big[\mathcal{B}_{1}(x_{1},\vartheta_{1});\ldots;\mathcal{B}_{N}(x_{N},\vartheta_{N})\big]$, condition (3.2c) is also met. The only remaining part is to show that $\gamma=1+\Gamma$ and $0<\gamma<1$. Given that $\xi^{\top}(-\Theta+\Psi):=\big[\xi_{1}\Gamma_{1};\dots;\xi_{N}\Gamma_{N}\big]^{\top}<0$, and $\max_{1\leq i\leq N}\Gamma_{i}<\Gamma<0$ with $\Gamma\in(-1,0)$, we have
@@
$=s+\big[\xi_{1}\Gamma_{1};\dots;\xi_{N}\Gamma_{N}\big]^{\top}\big[\mathcal{B}_{1}(x_{1},\vartheta_{1});\ldots;$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\mathcal{B}_{N}(x_{N},\vartheta_{N})\big]$
@@
Remark 4.3.
As the interconnection topology becomes denser, each subsystem is influenced by more neighbors, leading to a stronger aggregate disturbance $w_{i}$. To conservatively account for this in the analysis, the associated disturbance gain $\varphi_{i}$ may also increase, reflecting higher sensitivity to inter-subsystem coupling. Consequently, the compositional condition (4.1a), which relies on bounding this influence, becomes more difficult to satisfy, consistent with small-gain reasoning in stability analysis (see, e.g., (Dashkovskiy et al., 2007)).
As demonstrated in Theorem 4.2, under the compositionality conditions of (4.1a)-(4.1b), an A-CBC for the interconnected A-SHS can be compositionally constructed based on A-CSBCs of individual subsystems. In the next section, we provide the sufficient conditions for constructing the A-CSBC for the augmented SHS.
5 Construction of A-CSBC
In this section, we establish conditions on the SHS $\Phi_{i}$ that allow for the construction of an A-CSBC for $\mathcal{A}_{i}(\Phi_{i})$. To do so, we first define a notion of control barrier certificates (CBC) over original $\Phi_{i}$ without considering the sampling or jump constraints (Nejati et al., 2020; Anand et al., 2022). This intermediate formulation serves as a basis for systematically extending the certificates to the augmented setting to capture hybrid timing behavior.
@@ 5 Construction of A-CSBC
Given an SHS $\Phi_{i}$ with initial and unsafe sets $X_{0_{i}},X_{u_{i}}\subseteq X_{i}$, a function $\bar{\mathcal{B}}_{i}:X_{i}\rightarrow\mathbb{R}_{0}^{+}$ is called a control barrier certificate (CBC) for $\Phi_{i}$ if there exist $\gamma_{1_{i}}\in{\mathbb{R}},\bar{\kappa}_{i},\gamma_{2_{i}}\in{\mathbb{R}}^{+}$, $\bar{\mu}_{i},\bar{\beta}_{i},\bar{\varphi}_{i_{1}},\bar{\varphi}_{i_{2}},\bar{\eta}_{1_{i}},\bar{\eta}_{2_{i}}\in{\mathbb{R}}_{0}^{+}$, such that
$\bar{\mathcal{B}}_{i}(x_{i})\geq\bar{\kappa}_{i}\|x_{i}\|^{2},\quad\quad~~~~~\!\forall x_{i}\in X_{i},$ (5.1a)
$\bar{\mathcal{B}}_{i}(x_{i})\leq\bar{\mu}_{i},\quad\quad\quad\quad\quad\quad\!\!\forall x_{i}\in X_{0_{i}},$ (5.1b)
$\bar{\mathcal{B}}_{i}(x_{i})\geq\bar{\beta}_{i},\quad\quad\quad\quad\quad\quad\!\forall x_{i}\in X_{u_{i}},$ (5.1c)
@@
$\mathcal{L}\bar{\mathcal{B}}_{i}(x_{i})\leq-\gamma_{1_{i}}\bar{\mathcal{B}}_{i}(x_{i})+\bar{\varphi}_{i_{1}}\|w_{i}\|^{2}+\bar{\eta}_{1_{i}},$ (5.1d)
where $\mathcal{L}\bar{\mathcal{B}}_{i}(x_{i})$ is the infinitesimal generator acting on $\bar{\mathcal{B}}_{i}(x_{i})$, defined as (Oksendal, 2013)
$\mathcal{L}\bar{\mathcal{B}}_{i}(x_{i})$ $=\frac{\partial\bar{\mathcal{B}}_{i}(x_{i})}{\partial x_{i}}\!f_{1_{i}}\!(x_{i},\!\nu_{i},\!w_{i})\!$
$~~~+\frac{1}{2}\!\textsf{Tr}~\!(\sigma_{i}(x_{i})\sigma_{i}(x_{i})\!^{\top}\frac{\partial^{2}{\bar{\mathcal{B}}_{i}(x_{i})}}{\partial x_{i}\partial x_{j}})$
$~~~+\sum_{j=1}^{\textsf{r}_{i}}\lambda_{j_{i}}(\,\bar{\mathcal{B}}_{i}(x_{i}+\rho_{i}(x_{i})\textsf{e}_{j_{i}}^{\textsf{r}_{i}})-\bar{\mathcal{B}}_{i}(x_{i})),$
where $\lambda_{j}$ is the rate of Poisson processes and $\textsf{e}_{j}^{\textsf{r}}$ is an r-dimensional vector with $1$ on the $j$-th entry and 0 elsewhere;
@@
The above definition introduces a stochastic CBC $\bar{\mathcal{B}}_{i}$ for the SHS $\Phi_{i}$, analyzing the system behavior. For instance, with $\gamma_{1_{i}}>0$ and $0<\gamma_{2_{i}}<1$, it ensures that the expected certificate value decreases (or remains bounded) during both flow and jump phases (conditions (5.1d) and (5.1e), respectively). It is worthwhile noting that a thorough analysis is conducted for other bounds of $\gamma_{1_{i}}$ and $\gamma_{2_{i}}$ in Theorem 5.4 (cf. (5.8)). The terms $\bar{\varphi}_{i_{1}}$ and $\bar{\varphi}_{i_{2}}$ capture the effect of the adversarial input, enabling this certificate to quantify robustness against such disturbances, while for instance $\gamma_{1_{i}}>0$ and $0<\gamma_{2_{i}}<1$ characterize exponential decay or boundedness of the certificate in each mode.
Remark 5.2.
Definition 5.1 formulates CBCs tailored to SHSs that exhibit both continuous-time and discrete-time (reset) dynamics, as defined in Definition 2.1. In this setting, in the case that $\gamma_{1_{i}}>0$ and $0<\gamma_{2_{i}}<1$, condition (5.1d) ensures the desired decrease of the certificate along the continuous-time evolution, using the infinitesimal generator commonly employed in continuous-time stochastic analysis (Nejati et al., 2020). In contrast, condition (5.1e) governs the discrete-time updates and captures the expected decrease across discrete transitions, consistent with the discrete-time stochastic setting discussed in (Anand et al., 2022). Together, these conditions jointly ensure the safety of the SHS $\Phi_{i}$ under both modes of evolution.
Before offering the main result of this section, we also raise the following lemma to quantify an upper bound on the evolution of CBC $\bar{\mathcal{B}}_{i}$.
@@
$\mathds{E}$ $\Big[\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t_{k+1}^{-}))\,\big|\,x_{i},\nu_{i},w_{i}\Big]$
$\leq e^{-\gamma_{1_{i}}(t_{k+1}-t_{k})}\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t_{k})\!)$
$~~~+e^{-\gamma_{1_{i}}(t_{k+1}-t_{k})}(t_{k+1}\!-\!t_{k})(\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+\bar{\eta}_{1_{i}}).$ (5.2)
Proof. From Dynkin’s formula (Dynkin, 1965), we have:
$\mathds{E}$ $\Big[\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t_{k+1}^{-}))\,\big|\,x_{i},\nu_{i},w_{i}\Big]$
$=\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t_{k}))\!+\!\mathds{E}\Big[\int_{t_{k}}^{t_{k+1}}\!\!\!\!\!\!\!\mathcal{L}\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t))\mathsf{d}t\Big]\!.$
@@
$\mathds{E}$ $\Big[\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t_{k+1}^{-}))\,\big|\,x_{i},\nu_{i},w_{i}\Big]\leq\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t_{k}))$
$+\mathds{E}\Big[\int_{t_{k}}^{t_{k+1}}\big(-\gamma_{1_{i}}\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t))+\bar{\varphi}_{i_{1}}\|w_{i}\|^{2}+\bar{\eta}_{1_{i}}\big)\mathsf{d}t$
$=\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t_{k}))\!+\!(t_{k+1}\!-\!t_{k})(\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+\bar{\eta}_{1_{i}})$
$~~~-\int_{t_{k}}^{t_{k+1}}\!\!\gamma_{1_{i}}\mathds{E}\Big[\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t))\Big]\mathsf{d}t.$
By employing Gronwall inequality (Gronwall, 1919), we have
$\mathds{E}$ $\Big[\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t_{k+1}^{-}))\,\big|\,x_{i},\nu_{i},w_{i}\Big]$
$\leq e^{-\gamma_{1_{i}}(t_{k+1}-t_{k})}\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t_{k}))$
$+e^{-\gamma_{1_{i}}(t_{k+1}-t_{k})}(t_{k+1}\!-\!t_{k})(\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+\bar{\eta}_{1_{i}}),$
@@
Under Definition 5.1 and Lemma 5.3, we propose the next theorem to construct an A-CSBC for $\mathcal{A}_{i}(\Phi_{i})$ based on the CBC $\bar{\mathcal{B}}_{i}$.
@@
Given an SHS $\Phi_{i}=(X_{i},U_{i},\mathcal{U}_{i},W_{i},\mathcal{W}_{i},$ $\sigma_{i},\rho_{i},f_{1_{i}},\varsigma_{i},f_{2_{i}})$ with its associated A-SHS $\mathcal{A}_{i}(\Phi_{i})=(\mathbb{X}_{i},\mathbb{U}_{i},\mathds{W}_{i},\sigma_{i},\rho_{i},\varsigma_{i},\mathbb{F}_{i},\mathbb{Y}_{i},\mathbb{H}_{i})$, suppose $\bar{\mathcal{B}}_{i}$ is a CBC for $\Phi_{i}$. If
@@
$\varpi_{i}:=\left\{\hskip-1.42262pt\begin{array}[]{lr}1,\quad\quad\quad\,~~~~\text{if}~\gamma_{1_{i}}>0~\&~0<\gamma_{2_{i}}<1,\\ e^{\gamma_{1_{i}}\tau\alpha_{1_{i}}\vartheta_{i}},\quad\text{if}~\gamma_{1_{i}}>0~\&~\gamma_{2_{i}}\geq 1,\\ \gamma_{2_{i}}^{\frac{\vartheta_{i}}{\alpha_{2_{i}}}},\quad\quad~~~~\!\text{if}~\gamma_{1_{i}}\leq 0~\&~0<\gamma_{2_{i}}<1,\end{array}\right$. (5.8)
for some $0<\alpha_{1_{i}}<1$ and $\alpha_{2_{i}}>\varepsilon_{2_{i}}$.
@@
$\mathds{E}$ $\Big[\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(t_{k+1}^{-}))\,\big|\,x_{i},\nu_{i},w_{i}\Big]$
$\leq e^{-\gamma_{1_{i}}\tau}\big(\bar{\mathcal{B}}_{i}(x_{x_{i}\nu_{i}w_{i}}(0))\!+\!\tau(\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+\bar{\eta}_{1_{i}})\big),$
@@
      $=e^{-\gamma_{1_{i}}\tau}\mathcal{B}_{i}(x_{i},\!\vartheta_{i})\!+\!e^{-\gamma_{1_{i}}\!\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}$
      $~~~+\!e^{-\gamma_{1_{i}}\tau}\!\tau\bar{\eta}_{1_{i}}$
      $\leq e^{-\gamma_{1_{i}}\tau}\mathcal{B}_{i}(x_{i},\!\vartheta_{i})\!+\!e^{-\gamma_{1_{i}}\!\tau}\tau\bar{\varphi}_{i_{1}}(\|w_{i}\|_{\infty}^{2}\!+\!\|w_{i}\|^{2})$
      $~~~+e^{-\gamma_{1_{i}}\tau}\!\tau\bar{\eta}_{1_{i}}$
      $=e^{-\gamma_{1_{i}}\tau}\mathcal{B}_{i}(x_{i},\!\vartheta_{i})\!+\!e^{-\gamma_{1_{i}}\!\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|^{2}$
      $~~~+e^{-\gamma_{1_{i}}\!\tau}\tau\bar{\varphi}_{i_{1}}\|w_{i}\|_{\infty}^{2}+e^{-\gamma_{1_{i}}\tau}\!\tau\bar{\eta}_{1_{i}}.$
@@
      $\leq\gamma_{2_{i}}\bar{\mathcal{B}}_{i}(x_{i})\!+\!\bar{\varphi}_{i_{1}}\|w_{i}\|^{2}\!+\!\bar{\eta}_{2_{i}}$
      $=\gamma_{2_{i}}\mathcal{B}_{i}(x_{i},\vartheta_{i})\!+\!\bar{\varphi}_{i_{2}}\|w_{i}\|^{2}+\bar{\eta}_{2_{i}}.$
      By defining $\gamma_{i}=\max\{e^{-\gamma_{1_{i}}\tau},\gamma_{2_{i}}\}$, $\varphi_{i}=\max\{e^{-\gamma_{1_{i}}\tau}\tau\bar{\varphi}_{i_{1}},\bar{\varphi}_{i_{2}}\}$ and $\eta_{i}\!=\!\max\{e^{-\gamma_{1_{i}}\!\tau}\tau\bar{\varphi}_{i_{1}}\\ \|w_{i}\|_{\infty}^{2}+e^{-\gamma_{1_{i}}\tau}\!\tau\bar{\eta}_{1_{i}},\bar{\eta}_{2_{i}}\}$, we have
      $\mathds{E}\Big[\mathcal{B}_{i}(x_{i}^{\prime},\!\vartheta_{i}^{\prime})\big|x_{i},\!\nu_{i},\!w_{i},\!\vartheta_{i}\Big]$ $\leq\gamma_{i}\mathcal{B}_{i}(x_{i},\!\vartheta_{i})$
      $~~~+\varphi_{i}\|w_{i}\|^{2}\!+\!\eta_{i}.$